Imagine you are at a gas station filling your tank with gas. The function C(g) represents the cost C of filling up the gas tank with g gallons. Given the equation C(g)=3.50(g) Please help don't understand!!!!
What does the number 3.50 represent in this problem?

Found 2 solutions by solver91311, jim_thompson5910:
Answer by solver91311(24713)



Think about this for just a minute. You go to a store to buy something. You buy some number of this one thing -- like 3 loaves of bread or 5 t-shirts or 6 gallons of gasoline. How much money did you spend? The number of things times the price for one, right?

So if the total cost at the gas station is and represents the number of gallons of gas you bought, what must the 3.50 represent?


John


Answer by jim_thompson5910(35256)

The 3.50 is simply the price per gallon.


Ex: g=1 which means C(1) = 3.50(1) = 3.50 (ie 1 gallon of gas costs $3.50)

when g=2, C(2) = 3.50(2) = 7.00 (ie 2 gallons of gas cost $7.00)

when g=3, C(3) = 3.50(3) = 10.50 (ie 3 gallons of gas cost $10.50)

etc...